# A refinement of the Pontryagin-Thom theorem for unstable Thom spectra and its applications

### v1

@@ 1. Introduction
  • a stable vector bundle isomorphism $\Phi:TM\cong g^{*}V$ (i.e., there exists a sufficiently large positive integer $N$ and a bundle isomorphism $\Phi:TM\oplus\mathbb{R}^{N}\cong g^{V}\oplus\mathbb{R}^{N}$).
@@
$\epsilon:\Omega^{\textnormal{Spin, line}}_{n}\to\Omega^{O}_{n-1}\oplus\Omega^{SO}_{n}$
@@
As corollaries of our main theorems, we present two applications. The first application determines the non-negative integers $n$ for which there exists a pair $(M,\eta)$ consisting of an $(n+1)$-dimensional compact Spin manifold $M$ with boundary $S^{n}$ and a line subbundle $\eta$ of $TM$ such that $\eta$ is orthogonal to the boundary on $S^{n}$. The second application provides an alternative proof of the main result in [1]. In [1, Section 6], Question 2 asks for a description of the cobordism group defined in [1] in terms of the homotopy groups of a spectrum. Our work offers a solution to this question. Furthermore, using this description, we resolve the case $n\equiv 1\pmod{4}$, which was left open in [1].
@@ 3. The proof of theorem A
Since $M,g,\phi]=0\in\Omega(X;\xi-{\mathbb{R}}^{m})$, there exist a sufficiently large positive integer $N$, a sequence of $n$-dimensional closed manifolds $M=M_{0},M_{1},\ldots,M_{l}$ with $M_{l+1}=\emptyset$, continuous maps $g_{i}:M_{i}\to X$, and bundle isomorphisms $\Phi_{i}$ between $TM_{i}\oplus{\mathbb{R}}^{m+N}$ and $g_{i}^{*}\xi\oplus{\mathbb{R}}^{N}$. Note that we set $g_{0}=g$ and $\Phi_{0}=\phi\oplus{\mathbb{R}}^{N}$.
@@
Here, when $n$ is even, the image of the homomorphism $\pi_{n+1}(GL(n+2))\to\pi_{n+1}(S^{n+1})\cong{\mathbb{Z}}$ induced by the continuous map $GL(n+2)\to S^{n+1}$ (taking the $(n+2)$-th column) contains $\chi(S^{n+2})=2$ [11]. Consequently, $2[r\cdot c]=0\in\pi_{n}(GL^{+}(n+1))$ holds. Therefore, $a$ can be taken as either 0 or $1$, which completes the proof for this case.
@@
Suppose $M_{0}\sim M_{1}$ via a cobordism $N$ between them. We consider constructing a vector field on $N$ that points outwards along the boundaries $M_{0}$ and $M_{1}$. If such a vector field is nowhere zero, we can compose the trivialization of $TN$ with the reflection defined by this vector field. This allows us to swap the inward and outward directions of the $\oplus{\mathbb{R}}$ factor at both $M_{0}$ and $M_{1}$, thereby establishing symmetry. If the Euler characteristic of $N$ is non-zero, such a vector field may not exist globally; however, one can always construct such a vector field away from a finite set of points. By excising open neighborhoods of these points from $N$, we obtain the relation $M_{1}\sim M_{0}\cup\bigcup_{j=1}^{k}S^{n}$.
@@
We demonstrate this below. Let $X^{n+1}$ denote $S^{\frac{n+1}{2}}\times S^{\frac{n+1}{2}}$ if $n+1$ is divisible by 4, and $S^{\frac{n-1}{2}}\times S^{\frac{n+3}{2}}$ if n1n-1 is divisible by 4. Note that $X^{n+1}$ is a stably trivial, connected, closed $(n+1)$-dimensional manifold with Euler characteristic $4$ (this proof requires $n\neq 1$, as connectedness fails otherwise). Since a stably trivial closed manifold is almost parallelizable, $TX^{n+1}$ admits a trivialization on $X^{n}$ minus a single point.
@@
Applying the Poincaré-Hopf theorem to $X^{n+1}$, the image of this element under the map $S^{n},GL(n+1)]\to[S^{n},S^{n}]$ coincides with the Euler characteristic of $X^{n+1}$, which is $4$. By the lemma stated earlier, swapping the outward and inward directions of $\oplus{\mathbb{R}}$ transforms an element of degree $4$ into an element of degree 24=22-4=-2 in $S^{n},GL(n+1)]$. Consequently, an $S^{n}$ carrying an element of degree 2-2 (corresponding to $r\cdot c]$ since the degree of $r\cdot c]$ is $2$) is bounded by $X^{n+1}\setminus\mathrm{int}(D^{n+1})$.
@@
In particular, given a pair $(M,X)$ consisting of a closed oriented $3$-manifold $M$ and a nowhere-vanishing vector field $X$ on $M$, there exists a triple $(N,Y_{1},Y_{2})$ consisting of a compact oriented $4$-manifold NN and two linearly independent vector fields $Y_{1},Y_{2}$ on $N$ such that:
@@
This result is stronger than the fact that the $3$-dimensional oriented cobordism group is trivial; it means that we can impose additional structure on the bounding manifolds.
@@ 5. The proof of theorem C
Let $n$ be a non-negative integer and $m$ be a positive integer. Then $\pi_{-m}(BO(n)^{-L_{n}^{O}})$ is isomorphic to the (nm)(n-m)-dimensional cobordism group, $\Omega_{n-m}^{O}$. A similar theorem holds for $BSO$, $B\mathrm{Spin}$, and $B\mathrm{Pin}^{-}$. Here, for convenience, we define $\Omega_{n-m}^{O}=0$ when $n<m$.
@@
The second part of the proof follows the same logic as in Theorem C1, with $SO$ replaced by $\mathrm{Spin}$ and $O$ replaced by $\mathrm{Pin}^{-}$, up to the final step. It remains to determine the values of $n$ for which the map $p:\pi_{-1}(B\mathrm{Pin}^{-}(n+1)^{-(L^{\mathrm{Pin}^{-}}_{n+1}\oplus\mathrm{det}L^{\mathrm{Pin}^{-}}_{n+1})})\to\Omega^{\mathrm{Pin}^{-}}_{n}\oplus\Omega^{\mathrm{Spin}}_{n+1}\to{\mathbb{Z}}/2{\mathbb{Z}}$ is the zero map.
@@
Since $2e=0$ in odd dimensions (for the vector bundle, i.e., when $n+1$ is odd), $p$ is the zero map when $n$ is even. When $n$ is odd, considering the element $S^{n+1},{\mathbb{R}}^{2}\subset TS^{n+1}\oplus{\mathbb{R}}^{2}]$, its image under $p$ is the Euler characteristic of $S^{n+1}$, which is $2$. Thus, $2$ is contained in the image of $p$. It remains only to consider the parity of the values of $p$.
@@ 6. Application
Consider the following condition for a non-negative integer $n$: ”There exists a pair $(M,\eta)$ consisting of an $(n+1)$-dimensional compact $\mathrm{Spin}$ manifold $M$ with boundary $\partial M=S^{n}$ and a line subbundle $\eta$ of $TM$ such that $\eta|_{S^{n}}=(TS^{n})^{\perp}$ (i.e., $\eta$ is orthogonal to the boundary along $S^{n}$).”
@@
When $n$ is odd, according to [4], the given condition is equivalent to the Euler characteristic of $M$ being zero. Consequently, the proposition is equivalent to the existence of a closed $(n+1)$-dimensional $\mathrm{Spin}$ manifold with Euler characteristic $1$. As mentioned previously, such manifolds exist only when $n\equiv 7\pmod{8}$ (i.e., when the dimension $n+1$ is divisible by 8), in which case $M=\mathbb{HP}^{\frac{n+1}{4}}\setminus\mathrm{int}(D^{n+1})$ satisfies the condition.
@@
  • $\xi$ is a rank 2 oriented subbundle of $TW$.
@@
If this element is 0, then $M_{1}\sim M_{2}$ holds by the definition of the equivalence relation. If this element is the image of the non-zero element of ${\mathbb{Z}}/2{\mathbb{Z}}$, we can adjust it using the image of $1$ under the map ${\mathbb{Z}}/2{\mathbb{Z}}\to\pi_{1}(B\mathrm{Spin}(n-1)^{-L^{\mathrm{Spin}}_{n-1}})$. Specifically, we can replace $M_{1},\eta_{1}\oplus{\mathbb{R}}\subset TM_{1}\oplus{\mathbb{R}}]$ with another element $M_{1},\eta^{\prime}_{1}\oplus{\mathbb{R}}\subset TM_{1}\oplus{\mathbb{R}}]$ such that it maps to the same element $M_{1},{\mathbb{R}}\subset TM_{1}\oplus{\mathbb{R}}]$ in $\pi_{0}(B\mathrm{Spin}(n)^{-L^{\mathrm{Spin}}_{n}})$ (the base class remains unchanged), but differs by the non-trivial element in the fiber. This adjustment ensures that the classes map to the same element in $\pi_{-1}(B\mathrm{Spin}^{c}(n-1)^{-(L^{\mathrm{Spin}^{c}}_{n-1}\oplus\mathcal{O}_{n-1})})$. This proves sufficiency. ∎

### v2

@@ 1. Introduction
  • a stable vector bundle isomorphism $\Phi:TM\cong g^{*}V$ (i.e., there exists a sufficiently large positive integer NN and a bundle isomorphism $\Phi:TM\oplus\mathbb{R}^{N}\cong g^{V}\oplus\mathbb{R}^{N}$).
@@
$\epsilon:\Omega^{\textnormal{Spin, line}}_{n}\to\Omega^{\mathrm{Pin}^{-}}_{n-1}\oplus\Omega^{\mathrm{Spin}}_{n}$
@@
As corollaries of our main theorems, we present two applications. The first application determines the non-negative integers $n$ for which there exists a pair $(M,\eta)$ consisting of an $(n+1)$-dimensional compact Spin manifold $M$ with boundary $S^{n}$ and a line subbundle $\eta$ of TMTM such that $\eta$ is orthogonal to the boundary on $S^{n}$. The second application provides an alternative proof of the main result in [1]. In [1, Section 6], Question 2 asks for a description of the cobordism group defined in [1] in terms of the homotopy groups of a spectrum. Our work offers a solution to this question. Furthermore, using this description, we resolve the case $n\equiv 1\pmod{4}$, which was left open in [1].
@@ 3. The proof of theorem A
Since $M,g,\phi]=0\in\Omega(X;\xi-{\mathbb{R}}^{m})$, there exist a sufficiently large positive integer NN, a sequence of $n$-dimensional closed manifolds $M=M_{0},M_{1},\ldots,M_{l}$ with $M_{l+1}=\emptyset$, continuous maps $g_{i}:M_{i}\to X$, and bundle isomorphisms $\Phi_{i}$ between $TM_{i}\oplus{\mathbb{R}}^{m+N}$ and $g_{i}^{*}\xi\oplus{\mathbb{R}}^{N}$. Note that we set $g_{0}=g$ and $\Phi_{0}=\phi\oplus{\mathbb{R}}^{N}$.
@@
Here, when $n$ is even, the image of the homomorphism $\pi_{n+1}(GL(n+2))\to\pi_{n+1}(S^{n+1})\cong{\mathbb{Z}}$ induced by the continuous map $GL(n+2)\to S^{n+1}$ (taking the $(n+2)$-th column) contains $\chi(S^{n+2})=2$ [11]. Consequently, $2[r\cdot c]=0\in\pi_{n}(GL^{+}(n+1))$ holds. Therefore, $a$ can be taken as either 0 or 11, which completes the proof for this case.
@@
Suppose $M_{0}\sim M_{1}$ via a cobordism NN between them. We consider constructing a vector field on NN that points outwards along the boundaries $M_{0}$ and $M_{1}$. If such a vector field is nowhere zero, we can compose the trivialization of TNTN with the reflection defined by this vector field. This allows us to swap the inward and outward directions of the $\oplus{\mathbb{R}}$ factor at both $M_{0}$ and $M_{1}$, thereby establishing symmetry. If the Euler characteristic of NN is non-zero, such a vector field may not exist globally; however, one can always construct such a vector field away from a finite set of points. By excising open neighborhoods of these points from NN, we obtain the relation $M_{1}\sim M_{0}\cup\bigcup_{j=1}^{k}S^{n}$.
@@
We demonstrate this below. Let $X^{n+1}$ denote $S^{\frac{n+1}{2}}\times S^{\frac{n+1}{2}}$ if $n+1$ is divisible by 4, and $S^{\frac{n-1}{2}}\times S^{\frac{n+3}{2}}$ if n1n-1 is divisible by 4. Note that $X^{n+1}$ is a stably trivial, connected, closed $(n+1)$-dimensional manifold with Euler characteristic 44 (this proof requires $n\neq 1$, as connectedness fails otherwise). Since a stably trivial closed manifold is almost parallelizable, $TX^{n+1}$ admits a trivialization on $X^{n}$ minus a single point.
@@
Applying the Poincaré-Hopf theorem to $X^{n+1}$, the image of this element under the map $S^{n},GL(n+1)]\to[S^{n},S^{n}]$ coincides with the Euler characteristic of $X^{n+1}$, which is 44. By the lemma stated earlier, swapping the outward and inward directions of $\oplus{\mathbb{R}}$ transforms an element of degree 44 into an element of degree 24=22-4=-2 in $S^{n},GL(n+1)]$. Consequently, an $S^{n}$ carrying an element of degree 2-2 (corresponding to $r\cdot c]$ since the degree of $r\cdot c]$ is 22) is bounded by $X^{n+1}\setminus\mathrm{int}(D^{n+1})$.
@@
In particular, given a pair $(M,X)$ consisting of a closed oriented 33-manifold $M$ and a nowhere-vanishing vector field $X$ on $M$, there exists a triple $(N,Y_{1},Y_{2})$ consisting of a compact oriented 44-manifold NN and two linearly independent vector fields $Y_{1},Y_{2}$ on NN such that:
@@
This result is stronger than the fact that the 33-dimensional oriented cobordism group is trivial; it means that we can impose additional structure on the bounding manifolds.
@@ 5. The proof of theorem C
Let $n$ be a non-negative integer and $m$ be a positive integer. Then $\pi_{-m}(BO(n)^{-L_{n}^{O}})$ is isomorphic to the (nm)(n-m)-dimensional cobordism group, $\Omega_{n-m}^{O}$. A similar theorem holds for BSOBSO, $B\mathrm{Spin}$, and $B\mathrm{Pin}^{-}$. Here, for convenience, we define $\Omega_{n-m}^{O}=0$ when $n<m$.
@@
The second part of the proof follows the same logic as in Theorem C1, with SOSO replaced by $\mathrm{Spin}$ and $O$ replaced by $\mathrm{Pin}^{-}$, up to the final step. It remains to determine the values of $n$ for which the map $p:\pi_{-1}(B\mathrm{Pin}^{-}(n+1)^{-(L^{\mathrm{Pin}^{-}}_{n+1}\oplus\mathrm{det}L^{\mathrm{Pin}^{-}}_{n+1})})\to\Omega^{\mathrm{Pin}^{-}}_{n}\oplus\Omega^{\mathrm{Spin}}_{n+1}\to{\mathbb{Z}}/2{\mathbb{Z}}$ is the zero map.
@@
Since $2e=0$ in odd dimensions (for the vector bundle, i.e., when $n+1$ is odd), $p$ is the zero map when $n$ is even. When $n$ is odd, considering the element $S^{n+1},{\mathbb{R}}^{2}\subset TS^{n+1}\oplus{\mathbb{R}}^{2}]$, its image under $p$ is the Euler characteristic of $S^{n+1}$, which is 22. Thus, 22 is contained in the image of pp. It remains only to consider the parity of the values of pp.
@@ 6. Application
Consider the following condition for a non-negative integer $n$: ”There exists a pair $(M,\eta)$ consisting of an $(n+1)$-dimensional compact $\mathrm{Spin}$ manifold $M$ with boundary $\partial M=S^{n}$ and a line subbundle $\eta$ of TMTM such that $\eta|_{S^{n}}=(TS^{n})^{\perp}$ (i.e., $\eta$ is orthogonal to the boundary along $S^{n}$).”
@@
When $n$ is odd, according to [4], the given condition is equivalent to the Euler characteristic of $M$ being zero. Consequently, the proposition is equivalent to the existence of a closed $(n+1)$-dimensional $\mathrm{Spin}$ manifold with Euler characteristic 11. As mentioned previously, such manifolds exist only when $n\equiv 7\pmod{8}$ (i.e., when the dimension $n+1$ is divisible by 8), in which case $M=\mathbb{HP}^{\frac{n+1}{4}}\setminus\mathrm{int}(D^{n+1})$ satisfies the condition.
@@
  • $\xi$ is a rank 2 oriented subbundle of TWTW.
@@
If this element is 0, then $M_{1}\sim M_{2}$ holds by the definition of the equivalence relation. If this element is the image of the non-zero element of ${\mathbb{Z}}/2{\mathbb{Z}}$, we can adjust it using the image of 11 under the map ${\mathbb{Z}}/2{\mathbb{Z}}\to\pi_{1}(B\mathrm{Spin}(n-1)^{-L^{\mathrm{Spin}}_{n-1}})$. Specifically, we can replace $M_{1},\eta_{1}\oplus{\mathbb{R}}\subset TM_{1}\oplus{\mathbb{R}}]$ with another element $M_{1},\eta^{\prime}_{1}\oplus{\mathbb{R}}\subset TM_{1}\oplus{\mathbb{R}}]$ such that it maps to the same element $M_{1},{\mathbb{R}}\subset TM_{1}\oplus{\mathbb{R}}]$ in $\pi_{0}(B\mathrm{Spin}(n)^{-L^{\mathrm{Spin}}_{n}})$ (the base class remains unchanged), but differs by the non-trivial element in the fiber. This adjustment ensures that the classes map to the same element in $\pi_{-1}(B\mathrm{Spin}^{c}(n-1)^{-(L^{\mathrm{Spin}^{c}}_{n-1}\oplus\mathcal{O}_{n-1})})$. This proves sufficiency. ∎